### v1

@@ Abstract
We contribute an in-depth analysis of the workflows and tensions arising from generative AI (genAI) use in biomedical visualization (BioMedVis). Although genAI affords facile production of aesthetic visuals for biological and medical content, the architecture of these tools fundamentally limits the accuracy and trustworthiness of the depicted information, from imaginary (or fanciful) molecules to alien anatomy. Through 17 interviews with a diverse group of practitioners and researchers, we qualitatively analyze the concerns and values driving genAI (dis)use for the visual representation of spatially-oriented biomedical data. We find that BioMedVis experts, both in roles as developers and designers, use genAI tools at different stages of their daily workflows and hold attitudes ranging from enthusiastic adopters to skeptical avoiders of genAI. In contrasting the current use and perspectives on genAI observed in our study with predictions towards genAI in the visualization pipeline from prior work, our refocus the discussion of genAI’s effects on projects in visualization in the here and now with its respective opportunities and pitfalls for future visualization research. At a time when public trust in science is in jeopardy, we are reminded to first do no harm, not just in biomedical visualization but in science communication more broadly. Our observations reaffirm the necessity of human intervention for empathetic design and assessment of accurate scientific visuals. Supplemental study materials are available at https://osf.io/mbw86/.
@@ 1 BioMedVis Context
    Implementation/production: Design or program a novel visualization through mapping and rendering geometric/scalar data, often synthesizing multiple such sources, into the final image(s) [78, 26, 19, 6]. Biological and medical phenomena are often characterized by spatial data [60, 53, 35]. These spatial representations may be figurative, i.e., a close replication of the source material, and vary in style and degree of interactivity [78]. Concept-driven figurative visuals, e.g., “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationE and “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationF, are usually produced by designers or illustrators [29, 13] with the appropriate visual abstraction for the intended story when data are absent or lack the necessary fidelity. Data-driven figurative representations, as in “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationH, are typically produced by visualization developers [47, 26], although there is a natural midpoint where skill set and visual style meet, as in “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationG. Termed illustrative visualization [61, 68], this discipline draws inspiration from traditional illustration techniques to render scientific data. On the opposite end of the representational spectrum, visuals may be fully abstracted from their spatial context [41, 13]. Regardless of the representational approach, a developer or designer may collect and combine multiple, heterogeneous data sources to build the final visualization [34, 26].
@@ 2 Related Work
The academic and scientific research community at large is similarly grappling with the challenge of how to preserve scientific integrity with the advent of genAI tools in research [2, 8]. Summarizing a body of work, writing manuscripts with correct grammar, and citing relevant work for a topic are all proposed uses for genAI in different scientific communities, even within visualization. In contrast, our interview study focuses less on these text-to-text aspects in the modern scientific process to focus instead on the means of visualization design and development in a research or production context for topics in biology and medicine. Our emphasis on the need to preserve spatial relations in visuals more figurative than abstract introduce a different set of constraints than previously identified for genAI-assisted workflows.
@@
Within biomedical visualization, Kim et al. [37] experiment with OpenAI’s DALL-E 3 [56] to generate images of biomedical subject matter, including cell cultures, histological slides, and medical diagnostic imaging, e.g., X-rays. The authors laud the genAI tool’s ability to render “convincing” and “visually compelling” scientific visualizations, but note the gross inaccuracies, such as misrepresenting anatomy or including irrelevant subject matter, that limits its use in scientific visualization. Two other studies focus on medical illustrations, one related to aesthetic surgical applications [10], the other on corneal surgical procedures [50]. Both produce inaccurate results. Of note, the authors of these three studies appear to be primarily situated within medical and clinical contexts, and conclude that collaboration between medical experts and AI developers will be necessary for developing these tools. We emphasize the crucial and (currently) absent voice of the BioMedVis expert, whose experience and training in human-centred storytelling is an essential bridge between aesthetics and accuracy which is missing in these projects. Through this work, we profile the values and skills that BioMedVis experts bring to these conversations and surface potential opportunities for collaboration. We complement other voices from within the community that have explored genAI tools to support the narrative of a patient-centred medical visualization. For example, Mittenentzwei et al. [49] develop and evaluate a genAI-assisted workflow to produce photo-realistic characters representing patients in data-driven medical stories. This workflow is time-efficient and democratizes the otherwise specialized process of character modelling, rigging, and animation. Moreover, their evaluation showed a receptiveness to AI-generated images amongst viewers. In our study, we explore a broader range of illustrative biomedical visualization techniques used by BioMedVis experts that capture other potential workflows and uses of genAI.
@@ 3 Methods
We began this work out of a desire to understand how BioMedVis experts actually use genAI tools, which we see as a necessary step toward developing guidelines and resources to help unite our community in the ethical and practical use of this technology. As BioMedVis practitioners and visualization researchers ourselves, we have observed a sense of insulation from the existential concerns around genAI since OpenAI’s public release of ChatGPT in late 2022 [55]. While some BioMedVis experts spoke of genAI as a paradigm-shifting tool, we wanted to learn if and how colleagues use, and are affected by, these tools in their daily lives. We were interested to learn what, if any, of the predictions described in Schetinger et al. [62] had come to pass, particularly within the BioMedVis subcommunity, which works at the margins of spatial and figurative art in science. Through semi-structured interviews with BioMedVis professionals, we elicited personal perspectives and actual ways of working, with the goal of understanding how the field is responding to genAI and its impact on the daily production of visualizations for high-stakes health contexts. For study instruments and further details we refer to supplementary material (https://osf.io/mbw86/).
@@ 3.1 Participant recruitment
We used purposive and convenience sampling [69], which allowed us to identify prospective participants according to expertise and role(s) for a cross-sectional representation of our target community. We acknowledge convenience sampling because the majority of the Euro-North American community, where we focused our study, is connected through first- or second-degree relations with our author team. We aimed to capture a range of career stages (early, mid, and late or close to retirement); organization size (freelance or small studio to large academic institution); and seniority (junior artists to creative directors in industry; PhD/postdocs to principal investigators in research). We furthermore recruited participants from different disciplines within the community; we targeted recruitment efforts to medical illustrators and designers within the Association of Medical Illustrators [4] (from here on referred to as designers). We also recruited from the the Workshop for Visual Computing for Biology and Medicine (VCBM) [21] for participants with computer science or more technical backgrounds (hereon referred to as developers). Participant profiles are illustrated in Fig. 1.
@@ 3.2 Interviews
We obtained ethics approval through the University of Toronto Ethics Board. prior to data collection. To guide our interviews we developed a semi-structured interview template divided into four sections (see supplementary material for details). This structure afforded room for both organic discussion and deeper probing into topics of interest:
@@
    Focus on the nuts and bolts of the BioMedVis pipeline through a workflow activity, designed to elicit points of discussion regarding genAI in the pipeline, following a structure similar to related interview studies in visualization [62] and UX design [42]. Using a collaborative Miro board [48], we walk participants through the pipeline and ask them to describe their current, potential, and resistance to genAI use, and why (e.g., Is there something that stops you from using genAI tools?).
@@
    Explore ethical concerns (e.g., What are the ethical issues surrounding the use of AI in BioMedVis?). Close with participants’ thoughts and predictions about its future impact on the field.
@@ 3.3 Participants
We refer to participants with pseudonyms throughout this report to preserve anonymity and limit use of information that could lead to identifying individuals. Fig. 1 shows participants’ pseudonyms, role (designer/developer), gender, professional setting (academia/industry) and approximate experience level. Participants are arranged according to their attitudes towards genAI, i.e., from enthusiastic adopters to skeptical avoiders, based on their responses to attitude and value-based questions in their interviews. We discuss the resulting arrangement in detail in Sec. 6.1.
@@ 4.1 Stage 1: Research and Ideation
The research and ideation stage initiates all BioMedVis projects: identifying the problem that needs to be solved, uncovering the user’s or client needs, learning about the topic through a literature review (often about novel research), and ideation, i.e., coming up with possible solutions to the problem. For visual development, six designers use text-to-image prompting for inspiration. Abstract and other-worldly images help designers explore different visual styles to achieve a desired aesthetic (e.g. colour palettes), or atmosphere. These outputs, Jeff and Neil say, can help “jump-start” conversations with clients or other stakeholders early in the development pipeline, and are integrated with the BioMedVis expert’s trained scientific knowledge and visual storytelling techniques to create the final visuals. This aesthetic does not suit every BioMedVis workflow; Jules described genAI images as being “like something out of a dream or a movie” and that he “can’t think of an instance to use these tools for any sort of inspirational content.” Eight designers agree that the derivative “samey-ness of the images gets boring,” as Frank put it, is an issue; their work relies on clear messaging with a unique and recognizable visual style.
Text-to-text models are seen as helpful for six participants when conducting initial research on a (not necessarily novel) topic before validating the information themselves by going directly to relevant sources. Requiring a large volume of training data to perform well, participants felt ChatGPT typically provides reasonably reliable text information about well-established scientific content (e.g. what can be found in university-level science courses [64]), but is less suitable for summaries on emerging scientific research topics [64]. Irrelevant or hallucinated references remain a problem, as do invented new terms, such as the “green glowing protein.” Margaret, who came across this term in an experiment asking ChatGPT for a summary of the green fluorescent protein, is particularly skeptical of the utility of genAI tools in her work. Ray, Alan, and others describe the well-known issue of “the confident idiot,” i.e., a mismatch of self-awareness against real expertise, as particularly problematic for the biomedical field, which relies on accurate information for informed decision-making in health and clinical contexts.
@@ 4.2 Stage 2: Implementation / Production
Three developers note a potential for text-to-text models to assist with synthetic data generation and some analysis, corroborating findings from related studies [59, 31]. Kim sees the potential to improve clinical training by generating additional plausible variances of anatomic structures, i.e., fitting within the average range of anatomic variation, and could be observed in reality. Achieving this remains a challenge, however, while the availability of medical training data remains limited due to concerns regarding patient confidentiality and security. Alan notes that a feasible solution could be to develop and train a model on a local or sequestered environment, although the computing power required for this would likely not be able to compete with, e.g. ChatGPT-4. However, an adequate amount of training data for niche spatial biological and medical data is currently lacking, and presents a significant barrier to use in any environment, sequestered or otherwise, as Ray, Alan, and Lois observe. Two participants also expressed concerns over unknowingly generating the likeness of a real or living person in their work, or exposing a patient’s data. Either scenario is seen as an enormous breach of confidentiality and ethical violation in the field, issues which participants felt are less fraught in other domains.
Ten participants identify candidate tasks for genAI in automation, including generating boilerplate or segments of code, cleaning data, and debugging. This aligns with findings in related studies [62], and is the value proposition of tools like GitHub Copilot [28]. Depending on the participant’s pipeline and preferences, this may enable automating the “boring” stuff, i.e., tedious or not enjoyable, low-level, or human resource-intensive tasks, and is true for both designers and developers. Alan sees roughly a "90%-10% split" in his workflow with such automation, allowing him to focus his energy on more complex or creative problems, a sentiment echoed by most other participants. Neil sees the opportunity for genAI to speed up production altogether by: “reduc[ing] the amount of work that goes into making high-quality imagery, to reduce friction in the production process, and the goal is to eventually use AI to reduce the actual labour hours.”
However, skeptics like Lois and Margaret resist genAI tools, having already invested time to hone their process and skills. “I’m very good at coding in Python myself,” says Lois, so I don’t want to use ChatGPT for that.” The joy of such work was echoed by Octavia and Ray. From a design perspective, Frank comments, “I probably reach a flow state more often in my work than anything else […] when I’ve already figured out all the problems in this drawing and now I’m just painting, and I’m happy, I’m free. [That] could be the kind of mindless work that a robot could be doing. But then I’d probably be sad, because I like doing that, and my work would be less joyful in some ways.”
@@ 4.3 Stage 3: Dissemination
Delivery and/or dissemination of the project occurs at this stage. Visual artifacts may support clinical training or research, patient education, or public health initiatives. In this stage, participants find genAI tools most useful for short, text-based elements like figure captions and metadata or translating complex technical or medical jargon to an appropriate level depending on the communication goals. As Octavia expresses, “Can we actually use these kind of tools to translate […] from doctors to lay people, or from visualization experts to non visualization experts? For me, this is a very exciting topic.” Ray adds, “One of the things I’m impressed with is how good the translation works. That’s very helpful in terms of the accessibility of content […] for the broader public”.
@@
Overall, we see participants express a prevailing sense of “insulation” from the impacts of genAI given its obvious limitations, while others are more open to its transformative potential in the field, despite ethical concerns. Considering the pipeline as a whole, 13 participants already integrate genAI to some extent in their production workflows, but primarily for tasks auxiliary to what they generally see as the core of their work. Designers avoid genAI for the manual rendering of their work, while developers resist genAI for all but the most basic code snippets when writing code for image rendering. On the whole, the designer group express more favourable attitudes towards genAI tool adoption than developers, as Fig. 1 indicates. Whether developer or designer, we find participants’ attitudes regarding genAI cut across a wide spectrum, from deep skepticism (Margaret, Lois) and distrust regarding its use in BioMedVis (Ursula, Arthur) to fully embracing these tools throughout the pipeline (Jeff), as a means of supercharging creative output (Neil)–we map this landscape in Fig. 1. Others, like Octavia and Kim, are enthusiastic but are not directly involved in implementation work to be using genAI tools actively. The rest fall somewhere in the middle, cautiously optimistic and keeping themselves up-to-date on genAI development in order to best understand its use and limitations. On the whole, we observe that designers appear to be slightly stronger adopters than the developers.
@@
Stepping back from the BioMedVis pipeline, our reflexive thematic analysis [9] affords insights into BioMedVis experts’ broader perspectives regarding genAI use. These organize into three overarching themes and associated sub-themes (bolded in this section) which relate to accurate representation of complex spatial data, empathetic design, and balancing creative choices with accuracy and purpose.
@@
BioMedVis experts apply subject domain knowledge and 3D spatial understanding to correctly and meaningfully align visual mappings with 3D structures, as Alan, Ray, and Margaret explain. Anthony adds that the range of possible visual mappings is therefore limited as compared to other data types, and therefore limits genAI’s utility at this stage.
Furthermore, hyper-specialization within BioMedVis will continue to be seen as having a “protective” advantage, as Arthur and Jeff predict. The more complex, novel, or niche the subject matter, visual or technical approach, and/or modality, e.g. physical, interactive experiences (experiential design), with or without immersive tech, the less likely genAI is to serve as a collaborative partner. For example, VR-based surgical training, medical legal visualization that captures unique injuries or pathologies, and interactive exhibits, will likely remain the purview of those trained BioMedVis experts for a long time.
@@
BioMedVis experts leverage their knowledge of target audiences to create visualizations that support learning, insight-generation, and other goal-oriented tasks. Many participants stated that core to their work is close collaboration with relevant stakeholders—domain experts, clinicians, patient populations, students, or general audiences—in a project through design study [65] or contracted work. Fluent in not only the scientific content but also a range of communication strategies, the BioMedVis expert serves as a skilled translator and visual storyteller. As Jeff describes, “We are custom storytellers. It’s listening to the client, developing that relationship [and] discussing the problem as well as the solutions. And then there’s the craft of bringing that to life through visuals, and ensur[ing] [they] do the job[…].” Jules and Alan echo the emphasis on relational, “human connection” as one of the irreplaceable, as well as most rewarding, aspects of the work.
@@
Further exacerbating this issue is the (current) black-box nature of genAI models. Inaccurate or unreliable outputs, whether the anatomical visuals as exemplified in Fig “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationA–D or blocks of code, can mislead and diffuse responsibility. Participants questioned who should be held accountable in instances where genAI is used and lines of accountability blur. As Kim summarizes, “There should be someone who can explain the results. It is about trust, and […] about competence.” Ray expands on this notion, stating that “it should not be something that’s a black box and in the background. [You] need the corresponding user interfaces to establish what the visual mapping should be. […] That’s the danger, and it applies to all these stages. If you don’t know what’s going on and some magic happened, then potentially you’re not aware of what’s being shown, and you may miss something.” George adds that having “some tools to track the provenance of the data, like annotation tools or for data traceability” which would increase trust and explainability.
Designers similarly feel this sense of responsibility toward accurate and explainable visual design. Isaac notes, “we should stand behind the accuracy of it and we should be held responsible when we screw up.”. Ursula similarly asserts, “You need to have the highest standards […] for patient care. And patient education is 100%. There is no room for error in this.” Accountability is placed on the BioMedVis expert in medical legal visualization, as Arthur explains: “As an expert in the field, you may be called to testify as to, how do you know this re-creation your studio has made is accurate? What is your background? How do you know this much about the subject? I don’t think AI will ever have a place in that. Because how do you attest that some machine learning actually knew that particular person’s hepatic artery had a branch in a place where a surgeon might have cut it accidentally?”
@@
BioMedVis experts advocate for diverse, fair representation; genAI models cannot. Our participants express that part of the responsibility of the BioMedVis expert is to ensure, while working with respective stakeholders, that the visualizations they produce serve their intended purpose with sensitivity. Those involved in patient-centred projects raise repeated concerns about biases in training data that produce insensitive text and visual depictions of health or medical data, or limited opportunities for diverse representation. Martha asserts, “In medicine, most things in the past were done with a man as a standard human. So it’s also important to be mindful of diverse groups. It can be gender, but also ethnicity, and different regions. It’s about trust, this is a big topic for me.” Isaac further stresses the need for built-correctional safeguards against “unjust, historical mistreatment” of minority groups, to build trust and achieve fair representation.

### v2

@@ Abstract
We contribute an in-depth analysis of the workflows and tensions arising from generative AI (genAI) use in biomedical visualization (BioMedVis). Although genAI affords facile production of aesthetic visuals for biological and medical content, the architecture of these tools fundamentally limits the accuracy and trustworthiness of the depicted information, from imaginary (or fanciful) molecules to alien anatomy. Through 17 interviews with a diverse group of practitioners and researchers, we qualitatively analyze the concerns and values driving genAI (dis)use for the visual representation of spatially-oriented biomedical data. We find that BioMedVis experts, both in roles as developers and designers, use genAI tools at different stages of their daily workflows and hold attitudes ranging from enthusiastic adopters to skeptical avoiders of genAI. In contrasting the current use and perspectives on genAI observed in our study with predictions towards genAI in the visualization pipeline from prior work, we refocus the discussion of genAI’s effects on projects in visualization in the here and now with its respective opportunities and pitfalls for future visualization research. At a time when public trust in science is in jeopardy, we are reminded to first do no harm, not just in biomedical visualization but in science communication more broadly. Our observations reaffirm the necessity of human intervention for empathetic design and assessment of accurate scientific visuals. Supplemental study materials are available at https://osf.io/genaixbiomedvis/.
@@ 1 BioMedVis Context
    Implementation/production: Design or program a novel visualization through mapping and rendering geometric/scalar data, often synthesizing multiple such sources, into the final image(s) [78, 26, 19, 6]. Biological and medical phenomena are often characterized by spatial data [60, 53, 35]. These spatial representations may be figurative, i.e., a close replication of the source material, and vary in style and degree of interactivity [78]. Concept-driven figurative visuals, e.g., “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationE and “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationF, are usually produced by designers or illustrators [29, 13] with the appropriate visual abstraction for the intended story and audience, or when data are absent or lack the necessary fidelity. Data-driven figurative representations, as in “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationH, are typically produced by visualization developers [47, 26], although there is a natural midpoint where skill set and visual style meet, as in “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationG. Termed illustrative visualization [61, 68], this discipline draws inspiration from traditional illustration techniques to render scientific data. On the opposite end of the representational spectrum, visuals may be fully abstracted from their spatial context [41, 13]. Regardless of the representational approach, a developer or designer may collect and combine multiple, heterogeneous data sources to build the final visualization [34, 26].
@@ 2 Related Work
The academic and scientific research community at large is similarly grappling with the challenge of how to preserve scientific integrity with the advent of genAI tools in research [2, 8]. Summarizing a body of work, writing manuscripts with correct grammar, and citing relevant work for a topic are all proposed uses for genAI in different scientific communities, even within visualization. In contrast, our interview study focuses less on these text-to-text aspects in the modern scientific process and instead on the means of visualization design and development in a research or production context for topics in biology and medicine. Our emphasis on the need to preserve spatial relations in visuals more figurative than abstract introduces a different set of constraints than previously identified for genAI-assisted workflows.
@@
Within biomedical visualization, Kim et al. [37] experiment with OpenAI’s DALL-E 3 [56] to generate images of biomedical subject matter, including cell cultures, histological slides, and medical diagnostic imaging, e.g., X-rays. The authors laud the genAI tool’s ability to render “convincing” and “visually compelling” scientific visualizations, but note the gross inaccuracies, such as misrepresenting anatomy or including irrelevant subject matter, that limits its use in scientific visualization. Two other studies focus on medical illustrations, one related to aesthetic surgical applications [10], the other on corneal surgical procedures [50]. Both produce inaccurate results. Of note, the authors of these three studies appear to be primarily situated within medical and clinical contexts, and conclude that collaboration between medical experts and AI developers will be necessary for developing these tools. We emphasize the crucial and (currently) absent voice of the BioMedVis expert, whose experience and training in human-centred storytelling is an essential bridge between aesthetics and accuracy, which is missing in these projects. Through this work, we profile the values and skills that BioMedVis experts bring to these conversations and surface potential opportunities for collaboration. We complement other voices from within the community that have explored genAI tools to support the narrative of a patient-centred medical visualization. For example, Mittenentzwei et al. [49] develop and evaluate a genAI-assisted workflow to produce photo-realistic characters representing patients in data-driven medical stories. This workflow is time-efficient and democratizes the otherwise specialized process of character modelling, rigging, and animation. Moreover, their evaluation showed a receptiveness to AI-generated images amongst viewers. In our study, we explore a broader range of illustrative biomedical visualization techniques used by BioMedVis experts that capture other potential workflows and uses of genAI.
@@ 3 Methods
We began this work out of a desire to understand how BioMedVis experts actually use genAI tools, which we see as a necessary step toward developing guidelines and resources to help unite our community in the ethical and practical use of this technology. As BioMedVis practitioners and visualization researchers ourselves, we have observed a sense of insulation from the existential concerns around genAI since OpenAI’s public release of ChatGPT in late 2022 [55]. While some BioMedVis experts spoke of genAI as a paradigm-shifting tool, we wanted to learn if and how colleagues use, and are affected by, these tools in their daily lives. We were interested to learn what, if any, of the predictions described in Schetinger et al. [62] had come to pass, particularly within the BioMedVis subcommunity, which works at the margins of spatial and figurative art in science. Through semi-structured interviews with BioMedVis professionals, we elicited personal perspectives and actual ways of working, with the goal of understanding how the field is responding to genAI and its impact on the daily production of visualizations for high-stakes health contexts. For study instruments and further details, see supplementary material (https://osf.io/genaixbiomedvis).
@@ 3.1 Participant recruitment
We used purposive and convenience sampling [69], which allowed us to identify prospective participants according to expertise and role(s) for a cross-sectional representation of our target community. We acknowledge convenience sampling because the majority of the Euro-North American community, where we focused our study, is connected through first- or second-degree relations with our author team. We aimed to capture a range of career stages (early, mid, and late or close to retirement); organization size (freelance or small studio to large academic institution); and seniority (junior artists to creative directors in industry; PhD/postdocs to principal investigators in research). We furthermore recruited participants from different disciplines within the community; we targeted recruitment efforts to medical illustrators and designers within the Association of Medical Illustrators [4] (from here on referred to as designers). We also recruited from the the Workshop for Visual Computing for Biology and Medicine (VCBM) [21] for participants with computer science or more technical backgrounds (hereon referred to as developers). Participant profiles are illustrated in Fig.˜1.
@@ 3.2 Interviews
We obtained ethics approval through the University of Toronto Research Ethics Board prior to data collection. To guide our interviews we developed a semi-structured interview template divided into four sections (see supplementary material for details). This structure afforded room for both organic discussion and deeper probing into topics of interest:
@@
    Focus on the nuts and bolts of the BioMedVis pipeline through a workflow activity, designed to elicit points of discussion regarding genAI in the pipeline, following a structure similar to related interview studies in visualization [62] and UX design [42]. Using a collaborative Miro board [48], we walk participants through the pipeline and ask them to describe their current, potential, and resistance to genAI use, and why (e.g., Is there something that stops you from using genAI tools at x stage?).
@@
    Explore ethical concerns (e.g., What are the ethical issues surrounding genAI use in BioMedVis?). Close with participants’ thoughts and predictions about its future impact on the field.
@@ 3.3 Participants
We refer to participants with pseudonyms throughout this report to preserve anonymity and limit use of information that could lead to identifying individuals. Fig.˜1 shows participants’ pseudonyms, role (designer/developer), gender, professional setting (academia/industry) and approximate experience level. Participants are arranged according to their attitudes towards genAI, i.e., from enthusiastic adopters to skeptical avoiders, based on their responses to attitude and value-based questions in their interviews. We discuss the resulting arrangement in detail in Sec. 6.1.
@@ 4.1 Stage 1: Research and Ideation
The research and ideation stage initiates all BioMedVis projects: identifying the problem that needs to be solved, uncovering the user’s or client needs, learning about the topic through a literature review (often about novel research), and ideation, i.e., coming up with possible solutions to the problem. For visual development, six designers use text-to-image prompting for inspiration. Abstract and other-worldly images help designers explore different visual styles to achieve a desired aesthetic (e.g., colour palettes) or atmosphere. These outputs, Jeff and Neil say, can help “jump-start” conversations with clients or other stakeholders early in the development pipeline, and are integrated with the BioMedVis expert’s trained scientific knowledge and visual storytelling techniques to create the final visuals. This aesthetic does not suit every BioMedVis workflow; Jules described genAI images as being “like something out of a dream or a movie” and that he “can’t think of an instance to use these tools for any sort of inspirational content.” Eight designers agree that the derivative “samey-ness of the images gets boring,” as Frank put it, is an issue; their work relies on clear messaging with a unique and recognizable visual style.
Text-to-text models are seen as helpful for six participants when conducting initial research on a (not necessarily novel) topic before validating the information themselves by going directly to relevant sources. Requiring a large volume of training data to perform well, participants felt ChatGPT typically provides reasonably reliable text information about well-established scientific content (e.g., what can be found in university-level science courses [64]), but is less suitable for summaries on emerging scientific research topics [64]. Irrelevant or hallucinated references remain a problem, as do invented new terms, such as the “green glowing protein.” Margaret, who came across this term in an experiment asking ChatGPT for a summary of the green fluorescent protein, is particularly skeptical of the utility of genAI tools in her work. Ray, Alan, and others describe the well-known issue of “the confident idiot,” i.e., a mismatch of self-awareness against real expertise, as particularly problematic for the biomedical field, which relies on accurate information for informed decision-making in health and clinical contexts.
@@ 4.2 Stage 2: Implementation / Production
Three developers note a potential for text-to-text models to assist with synthetic data generation and some analysis, corroborating findings from related studies [59, 31]. Kim sees the potential to improve clinical training by generating additional plausible variances of anatomic structures, i.e., fitting within the average range of anatomic variation, and could be observed in reality. Achieving this remains a challenge, however, while the availability of medical training data remains limited due to concerns regarding patient confidentiality and security. Alan notes that a feasible solution could be to develop and train a model on a local or sequestered environment, although the computing power required for this would likely not be able to compete with, e.g. ChatGPT-4. However, an adequate amount of training data for niche spatial biological and medical data is currently lacking, and presents a significant barrier to use in any environment, sequestered or otherwise, as Ray, Alan, and Lois observe. Two participants also expressed concerns over unknowingly generating the likeness of a real or living person in their work, or exposing a patient’s data. Either scenario is seen as an enormous breach of confidentiality and ethical violation in the field, issues which participants feel are less fraught in other domains.
Ten participants identify candidate tasks for genAI in automation, including generating boilerplate or segments of code, cleaning data, and debugging. This aligns with findings in related studies [62], and is the value proposition of tools like GitHub Copilot [28]. Depending on the participant’s pipeline and preferences, this may enable automating the “boring” stuff, i.e., tedious or not enjoyable, low-level, or human resource-intensive tasks, and is true for both designers and developers. Alan sees roughly a “90%-10% split” in his workflow with such automation, allowing him to focus his energy on more complex or creative problems, a sentiment echoed by most other participants. Neil sees the opportunity for genAI to speed up production altogether by: “reduc[ing] the amount of work that goes into making high-quality imagery, to reduce friction in the production process, and the goal is to eventually use AI to reduce the actual labour hours.”
However, skeptics like Lois and Margaret resist genAI tools, having already invested time to hone their process and skills. “I’m very good at coding in Python myself,” says Lois, “so I don’t want to use ChatGPT for that.” The joy of such work was echoed by Octavia and Ray. From a design perspective, Frank comments, “I probably reach a flow state more often in my work than anything else […] when I’ve already figured out all the problems in this drawing and now I’m just painting, and I’m happy, I’m free. [That] could be the kind of mindless work that a robot could be doing. But then I’d probably be sad, because I like doing that, and my work would be less joyful in some ways.”
@@ 4.3 Stage 3: Dissemination
Delivery and/or dissemination of the project occurs at this stage. Visual artifacts may support clinical training or research, patient education, or public health initiatives. In this stage, participants find genAI tools most useful for short, text-based elements like figure captions and metadata or translating complex technical or medical jargon to an appropriate level depending on the communication goals. As Octavia expresses, “Can we actually use these kind of tools to translate […] from doctors to lay people, or from visualization experts to non visualization experts? For me, this is a very exciting topic.” Ray adds, “One of the things I’m impressed with is how good the translation works. That’s very helpful in terms of the accessibility of content […] for the broader public.”
@@
Overall, we see participants express a prevailing sense of “insulation” from the impacts of genAI given its obvious limitations, while others are more open to its transformative potential in the field, despite ethical concerns. Considering the pipeline as a whole, 13 participants already integrate genAI to some extent in their production workflows, but primarily for tasks auxiliary to what they generally see as the core of their work. Designers avoid genAI for the manual rendering of their work, while developers resist genAI for all but the most basic code snippets when writing code for image rendering. On the whole, the designers express more favourable attitudes towards genAI tool adoption than developers, as Fig.˜1 indicates. Whether developer or designer, we find participants’ attitudes regarding genAI cut across a wide spectrum, from deep skepticism (Margaret, Lois) and distrust regarding its use in BioMedVis (Ursula, Arthur) to fully embracing these tools throughout the pipeline (Jeff), as a means of supercharging creative output (Neil)–we map this landscape in Fig.˜1. Others, like Octavia and Kim, are enthusiastic but are not directly involved in implementation work to be using genAI tools actively. The rest fall somewhere in the middle, cautiously optimistic and keeping themselves up-to-date on genAI development in order to best understand its use and limitations. On the whole, we observe that designers appear to be slightly stronger adopters than the developers.
@@
Stepping back from the BioMedVis pipeline, our reflexive thematic analysis [9] affords insights into BioMedVis experts’ broader perspectives regarding genAI use. These perspectives organize into three overarching themes and associated sub-themes (bolded in this section) relating to accurate representation of complex spatial data, empathetic design, and balancing creative choices with accuracy and purpose.
@@
BioMedVis experts apply subject domain knowledge and 3D spatial understanding to correctly and meaningfully align visual mappings with 3D structures, as Alan, Ray, and Margaret explain. Anthony adds that the range of possible visual mappings is therefore limited as compared to other data types, and as such limits genAI’s utility at this stage.
Furthermore, hyper-specialization within BioMedVis will continue to be seen as having a “protective” advantage, as Arthur and Jeff predict. The more complex, novel, or niche the subject matter, visual or technical approach, and/or modality, e.g., physical, interactive experiences (experiential design), with or without immersive tech, the less likely genAI is to serve as a collaborative partner. For example, VR-based surgical training, medical legal visualization that captures unique injuries or pathologies, and interactive exhibits, will likely remain the purview of those trained BioMedVis experts for a long time.
@@
BioMedVis experts leverage their knowledge of target audiences to create visualizations that support learning, insight-generation, and other goal-oriented tasks. Many participants stated that core to their work is close collaboration with relevant stakeholders—domain experts, clinicians, patient populations, students, or general audiences—in a project through design study [65] or contracted work. Fluent in not only the scientific content but also a range of communication strategies, the BioMedVis expert serves as a skilled translator and visual storyteller. As Jeff describes, “We are custom storytellers. It’s listening to the client, developing that relationship [and] discussing the problem as well as the solutions. And then there’s the craft of bringing that to life through visuals, and ensur[ing] [they] do the job[…].” Jules and Alan echo the emphasis on relational “human connection” as one of the irreplaceable, as well as most rewarding, aspects of the work.
@@
Further exacerbating this issue is the (current) black-box nature of genAI models. Inaccurate or unreliable outputs, whether the anatomical visuals as exemplified in Fig “It looks sexy but it’s wrong.” Tensions in creativity and accuracy using genAI for biomedical visualizationA–D or blocks of code, can mislead and diffuse responsibility. Participants questioned who should be held accountable in instances where genAI is used and lines of accountability blur. As Kim summarizes, “There should be someone who can explain the results. It is about trust, and […] about competence.” Ray expands on this notion, stating that “it should not be something that’s a black box and in the background. [You] need the corresponding user interfaces to establish what the visual mapping should be. […] That’s the danger, and it applies to all these stages. If you don’t know what’s going on and some magic happened, then potentially you’re not aware of what’s being shown, and you may miss something.” George adds that having “some tools to track the provenance of the data, like annotation tools or for data traceability” would increase trust and explainability.
Designers similarly feel this sense of responsibility toward accurate and explainable visual design. Isaac notes, “We should stand behind the accuracy of it and we should be held responsible when we screw up.” Ursula similarly asserts, “You need to have the highest standards […] for patient care. And patient education is 100%. There is no room for error in this.” Accountability is placed on the BioMedVis expert in medical legal visualization, as Arthur explains: “As an expert in the field, you may be called to testify as to, how do you know this re-creation your studio has made is accurate? What is your background? How do you know this much about the subject? I don’t think AI will ever have a place in that. Because how do you attest that some machine learning actually knew that particular person’s hepatic artery had a branch in a place where a surgeon might have cut it accidentally?”
@@
BioMedVis experts advocate for diverse, fair representation; genAI models cannot. Our participants express that part of the responsibility of the BioMedVis expert is to ensure, while working with respective stakeholders, that the visualizations they produce serve their intended purpose with sensitivity. Those involved in patient-centred projects raise repeated concerns about biases in training data that produce insensitive text and visual depictions of health or medical data, or limited opportunities for diverse representation. Martha asserts, “In medicine, most things in the past were done with a man as a standard human. So it’s also important to be mindful of diverse groups. It can be gender, but also ethnicity, and different regions. It’s about trust, this is a big topic for me.” Isaac further stresses the need for built-in correctional safeguards against “unjust, historical mistreatment” of minority groups, to build trust and achieve fair representation.